OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

divide by -3 2x+6=4 2x+6=-4 solve for x

OpenStudy (anonymous):

subtract the 6 by 4 and -4 then divide by 2

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

okay so when I subtract 6 from four I get 2 and then I subtract -4 from 2?

OpenStudy (anonymous):

you have to isolate x first thing you do is isolate the absolute value which is 2x+6 so you divide -12 by -3 2 negatives make a positive now 2x+6=4 2x+6=-4

OpenStudy (anonymous):

subtract 6 into 4= -2 now 2x=-2 divide 2 into -2

OpenStudy (anonymous):

x = -5

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

now subtract 6 into -4 2x=-10 divide -10 by 2

OpenStudy (anonymous):

@Spidey dont just give answers explain how to do it please

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

okay so x=1 and x=5?

OpenStudy (anonymous):

dont forget your negatives :)

OpenStudy (anonymous):

Okay when you replace the variable x and it equal [-4] inside those line it comes out as + 4 and times negative three = -12

OpenStudy (anonymous):

Thank you so much @KlOwNlOvE and @Spidey

OpenStudy (anonymous):

Are you going to give me a medal?

OpenStudy (anonymous):

@Spidey this site is also not just for metals its to help people and receive help

OpenStudy (anonymous):

Alright don't get mad cause you didn't a medal first. But, let me stop being greedy

OpenStudy (anonymous):

im not mad i dont need medals i honestly think i have enough :) i got the title i wanted you only got one cuz you asked im sure of it